# Toward Agentic Software Engineering Beyond Code: Framing Vision, Values, and Vocabulary

### v1

@@ 4. Toward Agentic SE Values and Principles
  • Principle of Awareness and EducationAgentic SE researchers need to build awareness – While awareness of agentic AI is on the rise thanks to agentic AI systems, products, and news (including ethical incidents), awareness of agentic SE will need to be built through education and training within and across disciplines (in classrooms and workplaces) as well as with accreditation, certification, funding bodies, media, and society. Maintaining an agentic SE incident log will also help. Previous SE models have shown the importance of such efforts to the long-term future of those models and their associated communities.
  • Principle of Usable Guidelines and ToolsAgentic SE researchers need to study and enable the adoption of research knowledge and solutions into practice – Building knowledge and solutions is necessary but not sufficient to progress agentic SE. To achieve real-world impact, they need to be translated into usable guidelines and tools for practice. Mirroring and leveraging AI’s multi-modal generation, they can have more industry-friendly interfaces and mode of delivery (e.g., posters, executive summaries, graphical abstracts, short videos).
@@ 5. Toward Agentic SE Vocabulary
Well-defined terminology is a precursor to formalizing taxonomy (e.g. Sapkota et al. distinguish between AI agents and agentic AI and present a taxonomy of AI agent paradigms (Sapkota et al., 2025)) and ultimately, knowledge areas. While agentic SE is nascent, discrepancies in terminology usage are already apparent. Some are syntactic (e.g., agentic AI software engineer vs. AI software engineer vs. agentic software engineer). While others are reflective of deeper semantic and even philosophical issues. So we can lay solid foundations for agentic SE, here are some considerations for designing and using terminology for agentic SE:
@@
  • Coverageconceptual completeness and accuracy – to what extent does the term cover (nearly) all the goals and responsibilities of the agentic SE process area, actor, activity, or artefact it aims to support or automate? For example, a software engineer is more than a programmer, so can an agent that enacts programming tasks be termed a software engineer? Autonomous coding agent (Li et al., 2025) is a more accurate description.
@@
  • Consistencyuniformity of use, stability, adaptability – to what extent do different research groups around the world within SE use the terminology in the same way, both in form (syntax) and meaning (semantics)? how long do they remain stable? when and how do they adapt to changes over time?
@@ 6. A Deliberate and Desirable Paradigm Shift
Building on the foundational aspects of previous SE models (Section 2) and recent visions of agentic SE (e.g., (Hassan et al., 2025; Roychoudhury et al., 2025)), this paper proposes a ‘whole of processvision that accounts for human and agent actors, artefacts, and activities and other socio-technical concerns across all areas of agentic SE: traditional SE ones renewed for agents and multi-agent systems and new ones for human-agent ecosystems. It presents a preliminary set of guiding values and principles for agentic SE to be comprehensive, responsible, adaptive, foundational, and translational. It offers guidance on designing and using agentic SE vocabulary and welcomes community collaboration to progress these ideas further. Inevitably, proposed ideas end up being defined by practice (e.g., Waterfall’s sequential fate). We may not know what agentic SE truly looks like until it is studied empirically ‘in the wild’. Longitudinal studies to identify useful patterns in agentic repositories and forums and in-depth studies based in interviews, surveys, and experiments with agentic SE teams are needed.
Acknowledgment The author is grateful to Prof John Grundy and Hashini Gunatilake for their valuable feedback and inputs.

### v2

@@ 4. Toward Agentic SE Values and Principles
  • Principle of Awareness and Education Agentic SE needs to build awareness – While awareness of agentic AI is on the rise thanks to agentic AI systems, products, and news (including ethical incidents), awareness of agentic SE (as a discipline) will need to be built through education and training (in classrooms and workplaces) as well as with accreditation, certification, funding bodies, media, and society. Maintaining an agentic SE incident log will also help. Previous SE models have shown the importance of such efforts to the long-term future of those process models and their associated communities.
  • Principle of Usable Guidelines and Tools Agentic SE needs to study and enable the adoption of research knowledge and solutions into practice – Building knowledge and solutions is necessary but not sufficient to progress agentic SE. To achieve real-world impact, they need to be translated into actionable guidelines and tools for practice. Mirroring and leveraging AI’s multi-modal generation, they can have more industry-friendly interfaces and modes of delivery (e.g., posters, executive summaries, graphical abstracts, short videos).
@@ 5. Toward Agentic SE Vocabulary
Well-defined terminology is a precursor to formalizing taxonomy (e.g. Sapkota et al. distinguish between AI agents and agentic AI and present a taxonomy of AI agent paradigms (Sapkota et al., 2025)) and ultimately, knowledge areas. Although agentic SE is nascent, drifts and discrepancies in terminology usage are already apparent. Some are syntactic (e.g., agentic AI software engineer vs. AI software engineer vs. agentic software engineer). While others are reflective of deeper semantic and even philosophical issues. So we can lay solid foundations for agentic SE, here are some considerations for designing and using terminology for agentic SE:
@@
  • Coverageconceptual completeness and accuracy – to what extent does the term cover (nearly) all the goals and responsibilities of the agentic SE process area, actor, activity, or artefact it aims to support? For example, a software engineer is more than a coder, so an autonomous coding agent (Li et al., 2025) needs to be able to do more than coding to be termed an AI software engineer.
@@
  • Consistencyuniformity of use, stability, adaptability – to what extent do different technology and research groups in the SE world use the terminology in the same way, both in form (syntax) and meaning (semantics)? how long do they remain stable? how do they adapt to changes over time?
@@ 6. A Deliberate and Desirable Paradigm Shift
Building on the foundational aspects of previous SE models (Section 2) and recent visions of agentic SE (e.g., (Hassan et al., 2025; Roychoudhury et al., 2025)), this paper proposes a ‘whole of processvision that accounts for human and agent actors, artefacts, and activities and other socio-technical concerns across all areas of agentic SE: traditional SE ones renewed for agents and multi-agent systems and new ones for human-agent ecosystems. It presents a preliminary set of comprehensive, responsible, adaptive, foundational, and translational (CRAFT) values and principles for agentic SE. It offers guidance on designing and using agentic SE vocabulary, welcoming community collaboration to progress ideas. Inevitably, proposed ideas end-up being defined by practice (e.g., Waterfall’s sequential fate). We may not know what agentic SE truly looks like until it is studied empirically ‘in the wild’. Longitudinal studies to find useful patterns in agentic repositories and forums, and in-depth empirical studies based on interviews, surveys, and experiments with agentic SE teams, using inter- and intra-disciplinary approaches, are needed.
Acknowledgment Thanks to John Grundy, Hashini Gunatilake, Abhik Roychoudhury, and the reviewers for their valuable feedback.